# The Expressive Capacity of State Space Models: A Formal Language Perspective

### v2

@@ 3.4 Difficulty of PARITY
Conversely, we next establish a design choice in SSMs which limits their power in emulating finite-state-automata, establishing – in the finite-precision setting – an even stronger separation between existing SSM variants and traditional RNNs than the circuit complexity arguments in Merrill et al. [49]. The language PARITY, the language of bitstrings with an even number of ones, is recognized by a finite-state automaton with 2 states, and is straightforwardly encoded into a traditional RNN, even a linear one, with finite precision. PARITY, while in principle expressible for transformers [9], is empirically very hard to learn for transformers [6, 15], as it can provably only be represented in sharp minima [26]. A sufficiently general SSM could easily recognize it at $d=1$ by setting $h_{0}=1$, $A(e_{1})=-1$, $A(e_{0})=0$, $B\equiv 0$, so that the sign of the single entry of $h_{t}$ indicates the parity. Such an SSM would need to be non-time-invariant and require negative or complex gate values; i.e., satisfy neither time-invariant nor nonnegative. Such properties are indeed provably needed:
@@
The proof is in Appendix B.2; it examines inputs of the form $1^{N}$ and shows that the activations $z_{N}$ converge as $N\rightarrow\infty$, and thus cannot reliably encode the parity of $N$. As we show in Theorem 13, the same result holds even for SSMs evading nonnegative when they are time-invariant, at least when the coefficients have rational angles in the complex planes. All extant SSMs we surveyed (Appendix, Section A) satisfy either Nonnegative or time-invariant. Hypothetical SSMs evading both nonnegative and time-invariant would be strictly stronger and can represent not only PARITY, but all regular languages known to be in $\operatorname{TC}^{0}$ (Theorem 21).
3.5 Which Regular Languages can SSMs Model?
We now combine Theorems 1 and 2 to derive an exact characterizations of the regular languages that modern non-time-invariant SSMs such as Mamba can recognize or predictively model – the two notions turn out to coincide here – in the finite-precision setting. The key insight is that $\mathcal{L}_{FF}$ and PARITY are fundamental building blocks of two classes of regular languages: the star-free languages and their complement, the non-star-free languages [60, 46]:
@@ 3.5 Which Regular Languages can SSMs Model?
$\mathcal{L}_{FF}$ is star-free: there is a way to define it without Kleene star. PARITY is not star-free; any regular expression for it must involve the Kleene star. Even languages whose intuitive definition involves the Kleene star may turn out to be star-free. For example, $(01)^{*}$ has an expression avoiding the star. [^1] It is the union of $\epsilon$ with the intersection of $0\Sigma^{*}$, $\Sigma^{*}1$, with the complements of $\Sigma^{*}00\Sigma^{*}$ and $\Sigma^{*}11\Sigma^{*}$. Equivalently, a language is star-free if it can be defined logically using only first-order quantifiers and the order relation [60]. Also, $\mathcal{L}$ is non-star-free if and only if recognizing it involves counting modulo some $K$ [46]; PARITY being the simplest example. Modern non-time-invariant SSMs such as Mamba cannot perform modulo counting, but they can model all star-free languages:
@@
The class of SSMs satisfying nonnegative can predictively model a regular language $\mathcal{L}$ at finite precision if and only if $\mathcal{L}$ is star-free.
The proof in Appendix B.3 uses the Krohn-Rhodes theorem [39] to reduce all star-free languages to flip flop-like state tracking. Importantly, there are well-known constructive criteria for deciding whether a given automaton defines a star-free language [60]; hence, we have a decidable criterion for the finite-state tracking problems that such SSMs satisfying Nonnegative can solve.
This is much simpler than the situation for transformers, where an exact characterization of their power within the regular languages is complicated: Angluin et al. [3] show that a certain formal abstraction of transformers (masked unique hard attention) also recognizes exactly the star-free languages, but constructions of realistic transformers via Krohn-Rhodes in Liu et al. [44] do not inherently length generalize, and empirical research indicates difficulty in generalization even for certain simple star-free languages [6, 45]. Known length-generalizing constructions are limited to very simple subclasses such as the piecewise testable languages [71]. In contrast, for SSMs we have a single model per language, at finite precision and for arbitrarily long inputs. Thus, we expect that the SSM architecture confers an advantage in star-free state tracking problems when compared to transformers – a prediction we will find supported experimentally (Figure 3).
3.6 Unbounded Counting
Having characterized the regular languages modeled by SSMs, we now consider languages requiring unbounded counting [20], specifically, languages recognized by keeping track of one or more counters, where each character causes a specific increment or decrement to each counter [38, 27, 70, 40]. A prime example is the Dyck-1 language of well-formed strings over “(” and “)”; here a counter is incremented (decremented) whenever an opening (closing) bracket is encountered; a string is well-formed if and only if the counter is 0 at the end of the string. Some other relevant formal languages are Shuffle-Dyck-$k$ (the shuffles of multiple Dyck-1 languages), $a^{n}b^{n}$ – here, $a$ increments the counter and $b$ decrements it, and $a^{n}b^{n}c^{n}$ – here, there are two counters, one keeping track of $a^{n}b^{n}$ and one of $b^{n}c^{n}$ (See Appendix C.2). Such counter languages are fundamental as basic context-free (Dyck-1, $a^{n}b^{n}$) or context-sensitive (e.g., $a^{n}b^{n}c^{n}$) languages [30], and have been the subject of studies of both transformers [6] and RNNs [70]. Many such languages are modeled by SSMs:
@@ 3.6 Unbounded Counting
Consider the languages Dyck-1, Shuffle-Dyck-$k$, n-ary Boolean Expressions, $a^{n}b^{n}$, $a^{n}b^{n}c^{n}$, and $a^{n}b^{n}c^{n}d^{n}$, as defined in Appendix C.2. Then each of these is predictively modeled by an SSM.
The proof is in Appendix B.4. Intuitively, an SSM can directly implement the required counters by setting $A\equiv 1$ and by defining $B(e_{\sigma})$ to the be increment or decrement cased by $\sigma$. In modeling such languages, SSMs pattern with both transformers [6] and LSTMs [70].
@@ 3.7 Bounded Hierarchical Structure without Stacks
Figure 2: Test error on the validation set for $\mathcal{L}_{FF}$, in the setup of Liu et al. [45]. On both the in-distribution (green) and out-of-distribution (orange) setting, Mamba achieves near-zero test error, in agreement with Theorem 1, and avoiding the failures of transformers observed by Liu et al. [45].
It is generally agreed that hierarchical structure is a key aspect of language, and that comprehending language at a human-like level requires the computational ability to process such structures [12, 43, 18]. The fundamental data structure for processing hierarchical structure is the stack, where information is stored and removed as one traverses to higher and lower levels of hierarchical embedding [30]. We now show that SSMs’ counting ability can offer shortcuts even on such languages modeling hierarchical structure, eschewing the need for a stack.
A useful abstraction of hierarchical structure as relevant to natural language is the family of Dyck languages. The bounded-depth Dyck language $Dyck_{K,h}$ with $K$ types of parentheses and depth $h$ is the language of well-bracketed strings over $(_{1}$, $)_{1}$, $\dots$, $(_{K}$, $)_{K}$, such that the number of yet unclosed brackets never exceeds $h$ in any prefix [28, 74]. The classical unbounded-depth Dyck languages – the limit where $h\rightarrow\infty$ – play a fundamental role as the backbone of context-free languages via the Chomsky-Schützenberger theorem [12]. Bounding the depth reflects the fact that deep embedding is rare in natural language [36, 7]. Prior work has found that two-layer transformers [73] and traditional RNNs [28, 6] both model all $Dyck_{K,h}$ languages. The same turns out to hold for SSMs:
@@
The bounded-depth Dyck language $Dyck_{K,h}$ is predictively modeled by a two-layer SSM at finite precision, with $d=\mathcal{O}(h\log K)$.
The proof is in Appendix B.5. Intuitively, the first layer records the depth of each parenthesis using the ideas from Theorem 5, and the second layer keeps track of the last open bracket at each depth using Theorem 1. We note that, since $Dyck_{K,h}$ is star-free, Theorem 4 already guarantees the existence of representing SSMs, but the depth and width guaranteed by Theorem 6 is likely to be much better than what would be obtained by a black-box application of Theorem 4: As Hewitt et al. [28] show, $h\log K$ units is optimal up to constants and is attained by traditional RNNs and LSTMs. The SSM construction is very different from that of Hewitt et al. [28] for traditional RNNs (both simple RNNs and LSTMs), which directly simulates a stack. Our construction is similar to the transformer construction in Theorem 4.2 in [73], which however has to rely on specific positional encodings, unlike the SSM construction. This highlights that stacks are not the only way of simulating bounded hierarchical structure in recurrent architectures, and non-stack-based strategies can even attain the same optimal scaling of hidden units. Probing whether such stack-free shortcuts are learned by SSM-based LLMs is an exciting problem for future research.
Figure 3: Results on 27 formal languages, comparing our Mamba results (blue) with transformer results reported by [6] (orange), on in-distribution lengths (solid) and out-of-distribution lengths (dotted). As predicted by Theorem 4, Mamba performs strongly on star-free languages, and even shows perfect length generalization. Again as predicted by Theorem 4, it performs poorly on non-star-free languages. Results for transformers from [6] are mixed. Mamba also succeeds on learning the counter languages from Theorem 5, showing perfect accuracy at in-distribution lengths at in-distribution lengths, but length generalization lags behind transformers.
@@ 4 Experiments
We have derived a fine-grained theoretical characterization of expressiveness strengths and limitations of SSMs. We now show that our positive results can be instantiated and learned in a realistic SSM implementation, by evaluating a recent highly successful SSM, Mamba [23].
@@
We empirically instantiate Theorem 1 using the dataset released by [45], reflecting the language $\mathcal{L}_{FF}$ as defined in Section 3.3. Matching Figure 2 in Liu et al. [45], we evaluated both with in-distribution data, and with out-of-distribution data where the distance between read and write instructions tended to be larger. We evaluate for predicting the bits following $r$ instructions [^2] At all other positions, predictive modeling is trivial, as it only needs to consider the input symbol at that position., matching the “deterministic/clean” mode of Liu et al. [45], and considered predictions to be correct only if all predictions within a sequence were correct. We provide further details in Appendix D.2. A small one-layer [^3] Theorem 1 constructs a two-layer SSM. We hypothesize that Mamba uses its local convolution (Remark 18) to replace the lower layer from the construction in Theorem 1. Mamba model converged to 0 error in both validation sets after $\sim$ 1400 steps (Figure 2), compared to 500 steps for an LSTM reported by Liu et al. [45]. In contrast, Liu et al. [45] found transformers to keep making occasional mistakes despite training for 10K steps.
Test Suite from Bhattamishra et al. [6]
@@
To test our theoretical results on regular and counter languages (Theorems 2, 4, 5), we test Mamba on 27 formal languages, including 18 regular languages and 9 counter languages, based on a prior study comparing transformers and RNNs [6]. The regular languages include a popular benchmark [68] and various regular expressions; 11 are star-free. The counter languages include the languages covered by Theorem 5. Definitions of all languages are provided in Appendix C. We chose this test suite because it precisely covers Theorems 4 and 5, and we have proven (in)expressibility results for each language in the set.
Figure 4: As predicted by Theorem 6, Mamba with 2 layers can model bounded-depth Dyck languages. Results for test set with strings of length $700\leq n\leq 1400$.
@@
Following [6], we trained the model for predictive modeling, i.e., at each step, the model outputs a label indicating the set of possible next characters (4), including EOS when required. Following [6], we count the model’s response on an input string as correct if and only if predictive modelling was successful at all positions in the input. Such a evaluation setup makes random baselines fairly low, where a random predictor would have an accuracy exponentially small in $N$ in each of the $N$ positions. Training inputs have length in [1,50]; the model is then evaluated on held-out bins with length [1,50] and [51,100]. Further experimental details are in Appendix D.1.
@@
We show our Mamba results, together with Transformer results reported by Bhattamishra et al. [6], in Figure 3. LSTMs perform perfectly on all languages, and are thus not shown. In a striking confirmation of Theorem 4, Mamba learns all star-free languages with strong length generalization, and does poorly on all non-star-free languages. Transformers show somewhat more mixed performance, and do not always length-generalize even on the star-free languages. In agreement with Theorem 5, Mamba, similar to Transformers, learns the counter languages, though it struggles with length generalization when compared to transformers. The differences in Mamba’s performance between star-free and counter languages may stem from the fact that the construction for the former class (Theorem 4) is able to use finite precision and bounded state values at arbitrary input lengths, while the latter (Theorem 5) uses unbounded state values.
@@
To test Theorem 6, we recreate the experimental setup from Yao et al. [74]. Matching their Figure 4, we trained Mamba to predictively model $Dyck_{K,h}$ at $K=8$ and $h=10$. The training and the validation set contained samples of length $\leq 700$, while the test set contained samples of length $700\leq n\leq 1400$. Yao et al. [74] found both transformers and LSTMs achieved strong performance on this setup. We provide further details in Appendix D.3. Recall that Theorem 6 shows that two-layer SSMs can predictively model $Dyck_{K,h}$. We trained Mamba with 1 or 2 layers and varying dimensionality, finding that two layers can achieve essentially perfect performance across model sizes, even on the test set (Figure 4 and 5).

### v3

@@ 3.2 Difficulty of PARITY
PARITY, the language of bitstrings with an even number of ones, is recognized by a finite-state automaton with 2 states, and is straightforwardly encoded into a traditional RNN, even a linear one, with finite precision. It is in principle expressible for transformers (chiang2022overcoming), but is empirically hard for transformers to learn (bhattamishra2020ability; deletang2022neural), as it can provably only be represented in sharp minima (HahnRofin2024Sensitive). A sufficiently general SSM could easily recognize it at $d=1$ by setting $h_{0}=1$, $A(e_{1})=-1$, $A(e_{0})=0$, $B\equiv 0$, so that the sign of the single entry of $h_{t}$ indicates the parity (Figure 2). Such an SSM would need to be non-time-invariant and require negative or complex gate values; i.e., satisfy neither time-invariant nor nonnegative. Thus, these design choices necessitated by optimization, limit the power of an SSM in emulating finite-state-automata, establishing an even stronger separation between existing SSM variants and traditional RNNs than the circuit complexity arguments in merrill2024illusion
@@
The proof is in Appendix B.2; it examines inputs of the form $1^{N}$ and shows that the activations $z_{N}$ converge as $N\rightarrow\infty$, and thus cannot reliably encode the parity of $N$. It should be noted that we require the layer-wise operations used in the SSM to be either linear or based on the GLU or SwiGLU activation functions, as seen for instance in Mamba (Remark 15). As we show in Theorem 13, the same result holds even for SSMs evading nonnegative when they are time-invariant, at least when the coefficients have rational angles in the complex planes. All extant SSMs we surveyed (Appendix, Section A) satisfy either Nonnegative or time-invariant. Hypothetical SSMs evading both nonnegative and time-invariant would be strictly stronger and can represent not only PARITY, but all regular languages known to be in $\operatorname{TC}^{0}$ (Theorem 22).
Figure 3: (a) Construction for Flip-Flop (Theorem 1): The first layer stores instruction bits to the hidden state, while data bits are forwarded to the output. Hence, the output always contains both the latest instruction and the associated data bit. In the second layer, if the instruction bit is w, the corresponding data bit is written to the hidden state, else the old value persists. This allows the model to consistently output the correct data bit. (b) Construction for Dyck(K, h) (Theorem 6): The first layer tracks the depth by counting up for each opening bracket, and down for each closing bracket. The second layer builds on the Flip-Flop construction to find the last opening bracket at the current depth; the next symbol can be either the matching closing bracket or – if the maximum depth has not been reached – an arbitrary opening bracket.
3.3 Exact characterization of Regular Languages modeled by SSMs
We combine Theorems 1 and 2 to derive an exact characterizations of the regular languages that modern non-time-invariant SSMs such as Mamba can recognize or predictively model – the two notions coincide here – in the finite-precision setting. The key insight is that $\mathcal{L}_{FF}$ and PARITY are fundamental building blocks of two classes of regular languages: star-free languages and their complement, non-star-free languages (schutzenberger1965finite; mcnaughton1971counter):
@@ 3.3 Exact characterization of Regular Languages modeled by SSMs
$\mathcal{L}_{FF}$ is star-free: there is a way to define it without Kleene star. PARITY is not star-free; any regular expression for it must involve the Kleene star. Some languages that are intuitively defined with Kleene stars may still be star-free. [^2] For example, $(01)^{*}$ is star free. It is the union of $\epsilon$ with the intersection of $0\Sigma^{*}$, $\Sigma^{*}1$, with the complements of $\Sigma^{*}00\Sigma^{*}$ and $\Sigma^{*}11\Sigma^{*}$. A language is star-free if and only if it can be defined logically using only first-order quantifiers and the order relation (schutzenberger1965finite). Also, $\mathcal{L}$ is non-star-free if and only if recognizing it involves counting modulo some finite integer $K$ (mcnaughton1971counter); Modern non-time-invariant SSMs such as Mamba cannot perform modulo counting, but they can model all star-free languages:
@@
Let $\mathcal{L}$ be a regular language. The following are equivalent:
  1. 1.
    There is an SSM satisfying nonnegative that predictively models $\mathcal{L}$ at all input lengths, at finite precision
  2. 2.
    $\mathcal{L}$ is star-free.
The proof in Appendix B.3 uses the Krohn-Rhodes theorem (krohn1965algebraic) to reduce all star-free languages to flip flop-like state tracking. Importantly, there are well-known constructive criteria for deciding whether a given automaton defines a star-free language (schutzenberger1965finite); hence, we have a decidable criterion for the finite-state tracking problems that such SSMs satisfying Nonnegative can solve.
This is much simpler than the situation for transformers, where an exact characterization of their power within the regular languages is complicated: angluin2023masked show that a certain formal abstraction of transformers (masked unique hard attention) also recognizes exactly the star-free languages, but constructions of realistic transformers via Krohn-Rhodes in liu2022transformers do not inherently length generalize. Both theoretical (huang2024formal) and empirical research indicate difficulties in generalizing even for some simple star-free languages (bhattamishra2020ability; Liu2023FlipFlop). Known length-generalizing constructions are limited to very simple subclasses such as the piecewise testable languages (yang2024counting). In contrast, for SSMs we have a single model per language, at finite precision and for arbitrarily long inputs. Thus, we expect that the SSM architecture confers an advantage in star-free state tracking problems when compared to transformers – a prediction we will find supported experimentally (Figure 5).
3.4 SSMs can perform unbounded counting
Having characterized the regular languages modeled by SSMs, we now consider languages requiring unbounded counting (Fischer1968Counter), specifically, languages recognized by keeping track of one or more counters, where each character causes a specific increment or decrement to each counter (DBLP:conf/stacs/KrebsLL15; DBLP:conf/mfcs/HahnKLL15; weiss2018practical; kutrib2021input). A prime example is the Dyck-1 language of well-formed strings over “(” and “)”; here a counter is incremented (decremented) whenever an opening (closing) bracket is encountered; a string is well-formed if and only if the counter is 0 at the end of the string. Some other relevant formal languages are Shuffle-Dyck-$k$ (the shuffles of multiple Dyck-1 languages), $a^{n}b^{n}$ – here, $a$ increments the counter and $b$ decrements it, and $a^{n}b^{n}c^{n}$ – here, there are two counters, one keeping track of $a^{n}b^{n}$ and one of $b^{n}c^{n}$ (See Appendix C.2). Such counter languages are fundamental as basic context-free (Dyck-1, $a^{n}b^{n}$) or context-sensitive (e.g., $a^{n}b^{n}c^{n}$) languages (hopcroft2001introduction), and have been the subject of studies of both transformers (bhattamishra2020ability) and RNNs (weiss2018practical).
@@ 3.4 SSMs can perform unbounded counting
The languages Dyck-1, Shuffle-Dyck-$k$, n-ary Boolean Expressions, $a^{n}b^{n}$, $a^{n}b^{n}c^{n}$, and $a^{n}b^{n}c^{n}d^{n}$, (defined in Appendix C.2) can each be predictively modeled by an SSM.
The proof is in Appendix B.4. Intuitively (Figure 2), an SSM can directly implement the required counters by setting $A\equiv 1$ and by defining $B(e_{\sigma})$ to be the increment or decrpement cased by $\sigma$. In modeling such languages, SSMs pattern with both transformers (bhattamishra2020ability) and LSTMs (weiss2018practical).
@@ 3.5 Bounded Hierarchical Structure without Stacks
It is generally agreed that hierarchical structure is a key aspect of language, and comprehending language at a human-like level requires the computational ability to process such structures (chomsky1963algebraic; linzen2016assessing; everaert2015structures). The fundamental data structure for the same is a stack, where information is stored and removed as one traverses to higher and lower levels of hierarchical embedding (hopcroft2001introduction). We now show that SSMs’ counting ability can offer shortcuts on languages modeling hierarchical structure, eschewing the need for a stack.
A useful abstraction of hierarchical structure as relevant to natural language is the family of Dyck languages. The bounded-depth Dyck language $Dyck_{K,h}$ with $K$ types of parentheses and depth $h$ is the language of well-bracketed strings over $(_{1}$, $)_{1}$, $\dots$, $(_{K}$, $)_{K}$, such that the number of yet unclosed brackets never exceeds $h$ in any prefix (hewitt2020rnns; yao-etal-2021-self). The Chomsky-Schützenberger theorem (chomsky1963algebraic) asserts that any context-free language can be expressed as a homomorphic image of the intersection between a Dyck language and a regular language. Specifically, the Dyck language in question refers to the classical unbounded-depth Dyck language, where $h\rightarrow\infty$, underscoring its fundamental role as the structural backbone of context-free languages. Bounding the depth reflects the fact that deep embedding is rare in natural language (karlsson2007constraints; blasi2019distribution). Prior work has found that two-layer transformers (Yao_2021) and traditional RNNs (hewitt2020rnns; bhattamishra2020ability) both model all $Dyck_{K,h}$ languages. The same turns out to hold for SSMs:
@@
There is a two-layer SSM with $d=\mathcal{O}(h\log K)$ that predictively models $Dyck_{K,h}$ at all input lengths, at finite precision.
Figure 4: As predicted by Theorem 6, Mamba with 2 layers can model Dyck(K, h). Results for test set with strings of length $700\leq n\leq 1400$.
The proof is in Appendix B.5. Intuitively (Figure 3), the first layer records the depth of each parenthesis using the ideas from Theorem 5, and the second layer keeps track of the last open bracket at each depth using Theorem 1. We note that, since $Dyck_{K,h}$ is star-free, Theorem 4 already guarantees the existence of representing SSMs, but the depth and width guaranteed by Theorem 6 is likely to be much better than what would be obtained by a black-box application of Theorem 4: As hewitt2020rnns show, $h\log K$ units is optimal up to constants and is attained by traditional RNNs and LSTMs. The SSM construction is very different from that of hewitt2020rnns for traditional RNNs (both simple RNNs and LSTMs), which directly simulates a stack. Our construction is similar to the transformer construction in Theorem 4.2 in Yao_2021, which however has to rely on specific positional encodings, unlike the SSM construction. This highlights that stacks are not the only way of simulating bounded hierarchical structure in recurrent architectures, and non-stack-based strategies can even attain the same optimal scaling of hidden units. Probing whether such stack-free shortcuts are learned by SSM-based LLMs is an exciting problem for future research.
@@ 4 Experiments
We have derived a fine-grained theoretical characterization of expressiveness strengths and limitations of SSMs. We now show that our positive results can be instantiated and learned in a realistic SSM implementation, by evaluating a recent highly successful SSM, Mamba (Gu2023Mamba).
@@
We empirically instantiate Theorem 1 using the dataset of Liu2023FlipFlop, reflecting the language $\mathcal{L}_{FF}$ as defined in Section 3.1. Matching Figure 2 in Liu2023FlipFlop, we evaluated both on in-distribution data, and on out-of-distribution data where the distance between read and write instructions tended to be larger. We evaluate for predicting the bits following $r$ instructions [^3] Predictive modeling is trivial at other positions, as only the input symbols need to be considered there., matching the “deterministic/clean” mode of Liu2023FlipFlop, and considered predictions to be correct only if all predictions within a sequence were correct. (Further details in Appendix D.2). A small one-layer [^4] Theorem 1 constructs a two-layer SSM. We hypothesize that Mamba uses its local convolution (Remark 19) to replace the lower layer from the construction in Theorem 1. Mamba model converged to 0 error in both validation sets after $\sim$ 1400 steps (Figure 6), compared to 500 steps for an LSTM reported by Liu2023FlipFlop. In contrast, Liu2023FlipFlop found that transformers kept making occasional mistakes despite training for 10K steps.
Figure 5: Results on 27 formal languages, comparing our Mamba results (blue) with transformer results reported by bhattamishra2020ability (orange), on in-distribution lengths (solid) and out-of-distribution lengths (dotted). As predicted by Theorem 4, Mamba performs strongly on star-free languages, and even shows perfect length generalization. Again as predicted by Theorem 4, it performs poorly on non-star-free languages. Results for transformers from bhattamishra2020ability are mixed. Mamba also succeeds on learning the counter languages from Theorem 5, showing perfect accuracy at in-distribution lengths at in-distribution lengths, but length generalization lags behind transformers.
Figure 6: Test error on the validation set for $\mathcal{L}_{FF}$, following Liu2023FlipFlop. Mamba shows near-zero test error in both In- (green) and Out-of-distribution (orange) settings, consistent with Theorem 1, and avoids the failure seen in transformers (Liu2023FlipFlop)
Test Suite from bhattamishra2020ability
@@
To test our theoretical results on regular and counter languages (Theorems 2, 4, 5), we test Mamba on 27 formal languages, including 18 regular languages and 9 counter languages, based on a prior study comparing transformers and RNNs (bhattamishra2020ability). The regular languages include a popular benchmark (tomita1982dynamic) and various regular expressions; 11 are star-free. The counter languages include the languages covered by Theorem 5. (Definitions in Appendix C). We chose this test suite as it precisely covers Theorems 4 and 5, and we have proven (in)expressibility results for each language in the set.
@@
Following bhattamishra2020ability, we trained the model for predictive modeling, i.e., at each step, the model outputs a label indicating the set of possible next characters (3), including EOS when required. Following bhattamishra2020ability, we count the model’s response on an input string as correct if and only if predictive modelling was successful at all positions in the input. Such a evaluation setup makes random baselines low, where a random predictor would have an accuracy exponentially small in $N$ in each of the $N$ positions. Training inputs have length in [1,50]; the model is evaluated on held-out bins with length [1,50] and [51,100]. Further experimental details are in Appendix D.1.
@@
We show our Mamba results, together with Transformer results reported by bhattamishra2020ability, in Figure 5. LSTMs perform perfectly on all languages, and are thus not shown. In a striking confirmation of Theorem 4, Mamba learns all star-free languages with strong length generalization but performs poorly on non-star-free languages. Transformers show more mixed results, often failing to length-generalize even on star-free languages. Consistent with Theorem 5 , Mamba, like Transformers, learns counter languages but struggles more with length generalization. The differences in Mamba’s performance between star-free and counter languages may stem from the fact that the construction for the former class (Theorem 4) is able to use finite precision and bounded state values at arbitrary input lengths, while the latter (Theorem 5) uses unbounded state values.
@@
To test Theorem 6, we recreate the experimental setup from yao-etal-2021-self. Matching their Figure 4, we trained Mamba to predictively model $Dyck_{K,h}$ at $K=8$ and $h=10$. The training and the validation set contained samples of length $\leq 700$, while the test set contained samples of length $700\leq n\leq 1400$. yao-etal-2021-self found both transformers and LSTMs achieved strong performance on this setup. We provide further details in Appendix D.3. Recall that Theorem 6 shows that two-layer SSMs can predictively model $Dyck_{K,h}$. We trained Mamba with 1 or 2 layers and varying dimensionality, finding that two layers can achieve essentially perfect performance across model sizes, even on the test set (Figure 4 and 7).